A Complete Bibliography of Publications in the American Review of Mathematics and Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

13 October 2023 Version 1.05

Title word cross-reference

3 [Ali17, Ali18]. C [KG13]. C^{n} [Riz16]. C^{n+1} [Riz16]. e^{(aθ)} cos(bθ) [Kil19]. e^{(aθ)} sin(bθ) [Kil19]. e^{z} + 1/2z [Saj16]. F [Bar19, MMM22]. m [Ali19]. N [AK16, Ali19]. n = 3: 4 [Riz16]. P [ML17]. P^{0}(N) [Das13]. P^{d}(N) [Das13]. S [JA15, KG13]. ||a^{2}|| ≤ ||a||^{2} [MMM22]. ||fa|| = ||a|| [MMM22].

-component [Ali19]. -Dimensional [Ali17, Ali18, Ali19]. -Iterative [JA15]. -Space [AK16]. -Tests [KG13]. -Type [Bar19]. -Values [ML17].

Academic [Che15]. Achievement [Fri21, KG13]. Affecting [BOM17]. Against [Fod14]. Algebras [MMM22]. Algorithm [BGP15]. Allocation [Nma22, UAR16]. Almost [r15]. Alternate [La 16]. Analysis [Gup23, SSJ15, MHK^{+}16]. Anti [Reh16]. Anti-invariant [Reh16].

Application [Alh16, RRY23]. Applications [Kir14]. Applying [Kil19]. Approach [BOM17, Bar19, SSJ15]. Approximate [AM15]. Arbitrary [Li16]. Arguments [SK17]. Arrhythmia [Lin18]. Assessing [Bar19]. Assessment [TP15]. Asymptotic [xWX16]. Attribution [GS22b]. avoiding [BGP15].

B [AM15, SB16]. B-Spline [AM15]. B-Splines [SB16]. Bangkok [RRY23]. base [Ali19]. Based [Fri21, KG13, Lin18, MHK^{+}16]. Bases [Ali19]. Bayesian [Alh16, MHK^{+}16]. Beliefs [Fri21]. Better [Bar15]. Between [Lin17]. Biharmonic [Reh16]. Binary [BGP15]. Binet [Nai22]. Biochemical [TP15]. Black [GS22b]. Boundary [OKL14, SB16]. Branching [TT15]. Bridge [Che15].

C [Fri21]. Calculus [Tsu18]. Capacity [BOM17]. Case [Che15, JBA19]. Categories [Ced15]. Cement [BL15]. Central [GM15, MMM22]. Chains [GM15]. changing [MHK^{+}16]. Chemical [BL15]. Chlorophyll [TT15]. Chronic [MHK^{+}16]. Class [ASK13, Fin16, SB16]. Closed [CAR23]. Coefficient [Lin18]. College [Che15]. Comments [ML17]. Comparative [LE18, TP15]. Component [Alh16, Ali18, Ali19]. Composite [xWX16]. computer [Ali19]. conditions [Ali19, MHK^{+}16]. Confidence [MFH15, SI19]. Conservation [Fod14]. Contained [TT15]. Containing [MMM22]. Content [KG13]. Content-Based [KG13]. Continuous [Kum16]. Contractive [JA15]. Contractive-Like [JA15]. Control [Hou18]. Convergence [JA15]. Convex [AK16]. Cordial [MN16]. Correction [SB16]. Covariance [GM15]. CR [WC17]. CR-Submanifolds [WC17]. Crisp [Hag18]. Cuernavaca [JBA19].

Daily [RRY23]. Daniel [Fri21]. Data [JA15, MHK^{+}16]. Debt [GSVMEC14]. Defect [SB16]. Defined [Li16]. Defining [ML17]. Degree [RCK14, RCK16]. Dependence [JA15]. Detection [Lin18]. Determine [BOM17]. Deviating [SK17]. Difference [Tsu18]. Differentiable [AK16]. Differential [AM15, Dis15, Kum16, NF16, SK17]. Diffusion [xWX16]. Dimensional [Ali17, Ali18, Kum16, Ali19]. Dimensionless [II14]. Discrete [Tas21]. Discriminates [Riz16]. Discriminating [Lin17]. Disease [MHK^{+}16]. Distribution [ASK13, BN18, Ola15]. Distributions [Alk15]. Diversity [Hel22]. Divisor [MN16]. Do [Iza15]. Down [BGP15]. Driven [NF16]. duals [AR15]. Duplication [MN16]. Dynamic [OKL14].

Earthquakes [Lin17]. ECLIPSE [MHK^{+}16]. Economic [Ali17, Ali18, Ali19]. Economic-Mathematical [Ali17, Ali18, Ali19]. Educator [Fri21]. effective [Fri21]. Efficacy [MFH15]. Electron [Tas21]. Energy [Sak13, Tas21]. Equation [NF16]. Equations [AM15, Dis15, GSVMEC14, KN19, Kum16, OKL14, SK17]. Equivalence [AR15]. Equivalent [GSVMEC14]. Errors [Sak13]. Essential [Tas21].

Estimate [CAR23]. Estimates [Ced15]. Estimation [ASA22, xWX16]. Estimators [ASA22]. Euler [Kil19, Nai22]. Evaluation [Sak13]. even [Ali19]. Exchange [NUE16]. Exhaustive [BGP15]. Expansion [Li16]. Explosions [Lin17]. Extended [Alk15, Ola15]. Extreme [BN18].

Factors [Ali17, Ali18, BOM17]. Family [Alk15, Saj16]. Financial [NUE16]. Fit [Ced15]. Five [Tas21]. Flexible [Hou18]. Flux [Fod14]. following [JBA19]. Follows [Iza15]. Forecasts [Sak13]. Forms [Reh16]. Formula [Kil19]. Four [KG13]. Fourier [Li16]. Fractional [AM15, SK17]. frames [AR15]. Fredholm [Alt21]. Friedrichs [Fod14]. Function [AK16, Nai22, Saj16]. Functions [Das13, Li16, Ola15]. Fuzzy [Hag18].

Generalized [Alk15, Gup23, Ola15]. Generating [Das13]. Generation [BGP15]. germs [Riz16]. Godunov [Fod14]. Gompertz [BN18]. Goodness [Ced15]. Grade [KG13]. Graded [Ced15]. Graphs [MN16]. Gronwall [Kha15]. Groups [WY16]. GV [WY16]. GV-Semi [WY16].

Hartford [Che15]. Hausdorff [Dis15]. healing [BL15]. Heterogeneity [CAR23]. heterogeneous [MHK^{+}16]. Hierarchical [dA17]. High [KG13]. Highest [RRY23]. Hilbert [AR15, MMM22]. Hillyer [Che15]. Hubbard [Tas21]. Hypothesis [MFH15, ML17].

Idempotent [MMM22]. II [RCK16]. Implications [GSVMEC14]. Improved [SI19]. Improving [Che15, Fri21]. Impulses [Dis15]. Indicator [Fod14]. Inequalities [Kha15, Kir14]. Influence [Ali17, Ali18]. Inputs [Hag18]. Integral [Kha15, Kir14]. Integrate [Kil19]. Integro [AM15, Kum16]. Interval [SI19]. Intervals [Li16]. invariant [Reh16]. Inverse [BN18]. Investigation [Fod14]. Involving [Nai22]. Iraq [BOM17]. Itô [NF16]. Iterated [SB16]. Iterative [JA15].

Kaehlerian [WC17]. Kenmotsu [Reh16]. Kernel [dA17]. kernels [dA17].

Lattice [WY16]. Law [Iza15]. Laws [Fod14]. Lax [Fod14]. Lie [NF16]. Lifts [r15]. Like [JA15]. Limit [GM15]. Limited [Li16]. Linear [AM15, Ali17, Ali18, Bar19, Fin16, Gup23, LE18, Ali19]. Local [xWX16]. Locating [AK16]. Log [Iza15]. Logarithmic [Tsu18]. Logistic [Ola15]. Loop [GM15]. Loss [GS22b]. Lowest [RRY23].

Manifold [WC17]. Manifolds [Reh16]. map [Riz16]. Maps [Fin16]. Markov [GM15]. Master [RCK14, RCK16]. Mathematical [Ali17, Ali18, BL15, DCD15, Hel22, II14, Ali19]. Mathematics [RCK14, RCK16]. Mean [Tsu18]. Measures [Bar15, KG13]. Measuring [BOM17, Fri21]. Mechanism [BL15]. Memory [Iza15]. meta [MHK^{+}16]. meta-analysis [MHK^{+}16]. Method [AM15, Ali18, GS22a, Lin17]. Methods [Fod14]. Mexico [JBA19]. Mixed [Bar19]. Mixing [Hag18]. Mode [Ali18]. Model [BL15, Ced15, Che15, DCD15, GSVMEC14, GS22b, Hag18, Nma22, Sak13, Tas21, TT15, II14, xWX16, MHK^{+}16, Fri21]. model-based [MHK^{+}16]. Models [Ali17, Bar19, Hel22, Ali19]. Modern [RCK14, RCK16]. Modified [ASK13]. Morelos [JBA19]. most [Fri21]. Multi [Kum16]. Multiple [Bar19, MHK^{+}16]. Multitype [TT15]. multivariante [Ali19]. Multivariate [Ali18].

Negatively [Ola15]. Networks [TP15]. Nigerian [NUE16]. Non [AK16]. Non-Differentiable [AK16]. Nonlinear [Fod14, Kha15, SB16]. nonzero [MMM22]. Normal [WC17]. Notes [r15]. Null [ML17]. Number [Ced15]. Numbers [Iza15].

Observation [MFH15]. Obstructive [MHK^{+}16]. One [Saj16]. Operator [Tas21]. Operators [GM15, JA15]. Optimal [UAR16]. Optimization [NUE16]. Optimizer [AK16]. Option [Bar15]. Options [GS22b]. Orbital [Dis15]. Ordinal [Gup23]. Oscillation [SK17]. Overdue [MFH15].

Paracontact [r15]. Parameter [Saj16]. Parametric [SSJ15]. Pareto [BN18]. Partial [SK17]. Patterns [CAR23]. Pencils [Alt21]. Perception [Fri21]. Performance [Che15]. Periodic [Iza15]. Phytoplankton [TT15]. Picard [JA15]. Piecewise [Ali17, Ali18, Fin16, Ali19]. Piecewise-Linear [Ali18, Ali19]. Plate [Hou18]. PLS [Hös23]. Point [SB16]. Poisson [NF16]. Polynomial [Li16, Nma22]. Polynomial-Time [Nma22]. Population [ASA22, CAR23, TT15]. Portfolio [NUE16]. Positive [OKL14]. Potential [KN19]. Power [Alk15]. Practical [GSVMEC14]. Pre [MMM22]. Pre-Hilbert [MMM22]. Precision [Ced15]. Prediction [Ali18, Ali19]. Principal [Alh16]. Probability [Bar15, MHK^{+}16]. Problem [DCD15]. Problems [SB16]. Process [Ali18, NF16, TT15]. Profit [GS22b]. Program [Che15]. Programs [RCK14, RCK16]. Properties [ASK13, Alt21, xWX16, WY16]. Psychometric [Ced15]. Pulmonary [MHK^{+}16].

Quantile [BOM17, xWX16]. Quartet [Tas21]. Quasi [Ola15, WC17]. Quasi-Range [Ola15]. Quenching [KN19]. Quest [Bar15].

Random [Iza15]. Randomness [Bar19, Iza15]. Range [Ola15]. Reaction [BL15]. Redundancy [Nma22]. Regard [Ali17, Ali18]. Regression [Alh16, BOM17, Gup23, Hag18, Hös23, LE18, xWX16]. Reliability [UAR16]. Representations [Nai22]. Representative [GS22a]. Response [Ced15]. Restructuring [GSVMEC14]. Result [JA15]. Retarded [Kha15]. Ridge [Hös23]. RKDG [Fod14]. Robust [LE18]. Rule [JBA19]. Rule-following [JBA19].

SARFIMA [Sak13]. Scalar [Fod14]. Scale [Ced15, Kum16]. Scenario [GS22a]. Scheme [JA15]. Scholes [GS22b]. School [Fri21]. School-Based [Fri21]. Schools [KG13]. Science [Hel22]. SDFA [Lin17]. see [Fri21]. Self [BL15]. Self-healing [BL15]. Semi [WY16]. Semilinear [KN19]. Sequences [BGP15, Iza15]. Series [Alk15, Li16, RRY23]. Services [NUE16]. Sextet [Tas21]. Shearlet [Kum16]. Shrinkage [Kum16, Lin18]. Similarity [Hel22]. Simulation [TP15, Ali19]. Singular [Saj16]. Size [ASA22, CAR23]. Skewed [Ola15]. Smoothness [Fod14]. software [Ali19]. Solution [AM15]. Solutions [Dis15, OKL14, SK17]. Some [KN19, Nai22, r15]. Space [AK16, Ali17, Ali18, Reh16, Ali19]. spaces [AR15]. Spectra [Tas21]. Spectrum [Tas21]. Spline [AM15]. Splines [SB16]. Stability [Dis15]. States [Tas21]. Statistical [MFH15, SSJ15]. Stochastic [NF16]. Stock [NUE16]. Stocks [NUE16]. Strategies [Fri21]. Strength [Fri21]. Strength/Beliefs [Fri21]. Strongly [SB16]. Structural [Hou18]. Structure [Dis15, Tas21]. Structures [r15]. Student [Che15, Fri21]. Study [Che15, JBA19, LE18]. Sub [dA17, Reh16]. Sub-kernels [dA17]. Submanifolds [WC17]. such [MMM22]. Suggestions [ML17]. Summer [Che15]. Sums [Nai22]. Survey [RCK14, RCK16]. Switching [MN16]. Symmetry [NF16]. System [AM15]. Systems [Tas21].

TELBS [LE18]. Temperatures [RRY23]. Test [La 16, MFH15]. Testing [Bar19, MFH15]. Testlets [SSJ15]. Tests [KG13, ML17]. Thailand [RRY23]. their [AR15]. Theoretical [GSVMEC14]. Theory [Bar15, BN18, GM15]. Therein [TT15]. Three [Ali18]. Three-Component [Ali18]. Time [KN19, Nma22, RRY23, xWX16, MHK^{+}16]. Time-Varying [xWX16]. Tobit [Alh16]. Tools [TP15]. Total [ASA22]. Total-Utilizing [ASA22]. Tournament [DCD15]. Traffic [JBA19]. Transform [Kum16]. Traveling [DCD15]. Two [SB16]. Two-Point [SB16]. Type [Bar19, Kha15].

U.S. [RCK14, RCK16]. Unaccounted [Ali17, Ali18]. uncertainty [Ali19]. University [Che15]. Up-Down [BGP15]. Using [AM15, BOM17, Fod14, Fri21, JA15, ML17]. Utilizing [ASA22].

Value [BN18, OKL14, SB16]. Values [ML17, Saj16]. Variable [Dis15, Hou18]. Variances [La 16]. Varying [Ced15, CAR23, xWX16]. Vector [Ali17, Ali18, SK17, Ali19]. Volterra [AM15].

Wavelet [Kum16, Lin18]. Weibull [ASK13, Alk15]. Weighted [ASK13]. which [Fri21]. Words [BGP15].

References

Achi:2016:LON

[AK16] Godswill U. Achi and Agomoh Ugochukwu K. Locating the optimizer of a non-differentiable convex function in N-space. American Review of Mathematics and Statistics, 4(1):61–72, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Alhusseini:2016:BTP

[Alh16] Fadel Hamid Hadi Alhusseini. Bayesian tobit principal component regression with application. American Review of Mathematics and Statistics, 4(2):63–73, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Aliyev:2017:PLE

[Ali17] Azad Gabiloglu Aliyev. Piecewise linear economic-mathematical models with regard to unaccounted factors influence in 3dimensional vector space. American Review of Mathematics and Statistics, 5(1):7–28, June 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Aliyev:2018:TCP

[Ali18] Azad Gabiloglu Aliyev. Three-component piecewise-linear economic-mathematical mode and method of multivariate prediction of economic process with regard to unaccounted factors influence in 3-dimensional vector space. American Review of Mathematics and Statistics, 6(1):37–62, June 2018. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Aliyev:2019:BSC

[Ali19] Azad Gabiloglu Aliyev. Bases of software for computer simulation and multivariante prediction of economic even at uncertainty conditions on the base of n-component piecewise-linear economicmathematical models in m-dimensional vector space. American Review of Mathematics and Statistics, 7(1):12–36, June 2019. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Alkarni:2015:GEW

[Alk15] Said H. Alkarni. Generalized extended Weibull power series family of distributions. American Review of Mathematics and Statistics, 3(2):53–68, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Altaweel:2021:FPP

[Alt21] Nifeen Altaweel. Fredholm properties for pencils. American Review of Mathematics and Statistics, 9(1):1–11, June 2021. CO-DEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Al-Marashi:2015:ASS

[AM15] Adel A. Al-Marashi. Approximate solution of the system of linear fractional integro–differential equations of Volterra using B-spline method. American Review of Mathematics and Statistics, 3(2): 39–47, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Alijani:2015:EFT

[AR15] Azadeh Alijani and Zohreh Rahbani. Equivalence frames and their duals in Hilbert spaces. American Review of Mathematics and Statistics, 3(2):61–69, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Al-Saleh:2022:EPT

[ASA22] Mohammad Fraiwan Al-Saleh and Ma’aly Salah Ababneh. Estimation of the population total-utilizing estimators of the population size. American Review of Mathematics and Statistics, 10 (1):17–33, June 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Aleem:2013:CMW

[ASK13] M. Aleem, M. Sufyan, and N. S. Khan. A class of modified weighted Weibull distribution and its properties. American Review of Mathematics and Statistics, 1(1):29–37, December 2013. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (elec-

tronic). URL http://armsnet.info//journals/arms/Vol_1_
No_1_December_2013/3.pdf.

Barnard:2015:OPT

[Bar15] John J Barnard. Option probability theory: a quest for better measures. American Review of Mathematics and Statistics, 3 (1):61–69, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Barnabani:2019:TMT

[Bar19] Marco Barnabani. An F -type multiple testing approach for assessing randomness of linear mixed models. American Review of Mathematics and Statistics, 7(2):15–28, December 2019. CO-DEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Bilotta:2015:AEG

[BGP15] Stefano Bilotta, Elisabetta Grazzini, and Elisa Pergola. An algorithm for the exhaustive generation of binary words avoiding up-down sequences. American Review of Mathematics and Statistics, 3(1):9–21, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Bell:2015:MMC

[BL15] Joseph E. Bell and Dr. Tuwaner Lamar. A mathematical model of a chemical reaction mechanism for self-healing cement. American Review of Mathematics and Statistics, 3(1):1–8, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Bashir:2018:GIP

[BN18] Shakila Bashir and Itrat Batool Naqvi. The Gompertz inverse Pareto distribution and extreme value theory. American Review of Mathematics and Statistics, 6(2):30–37, December 2018. CO-DEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Bager:2017:UAQ

[BOM17] Ali Sadig Mohommed Bager, Meshal Harbi Odah, and Bahr Kadhim Mohammed. Using approach quantile regression to determine the factors affecting measuring capacity in Iraq. American Review of Mathematics and Statistics, 5(1):35–44, June 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Chen:2023:ESC

[CAR23] Lihua Chen, Kara Athey, and Nat Remmers. Estimate the size of a closed population with varying heterogeneity patterns. American Review of Mathematics and Statistics, 11(1):22–28, June 2023. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Cedeno:2015:GFP

[Ced15] Yuliana Mora Cedeño. Goodness of fit and precision of the graded response model estimates of a psychometric scale at varying number of categories. American Review of Mathematics and Statistics, 3(2):25–29, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Cheser:2015:UHH

[Che15] Elena A. Cheser. The University of Hartford, Hillyer College Summer Bridge Program: a model and case study for improving student academic performance. American Review of Mathematics and Statistics, 3(1):27–33, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Casanova-del-Ange:2017:HKS

[dA17] Francisco Casanova del Ange. Hierarchical kernel and subkernels. American Review of Mathematics and Statistics, 5(1): 29–34, June 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Das:2013:GF

[Das13] Sabuj Das. Generating functions for P^{0}(N) and P^{d}(N). American Review of Mathematics and Statistics, 1(1):38–40, December 2013. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (elec-

tronic). URL http://armsnet.info//journals/arms/Vol_1_
No_1_December_2013/4.pdf.

Dorrepaal:2015:MMT

[DCD15] J. Mark Dorrepaal, George Chackman, and Brian Dorrepaal. A mathematical model for the traveling tournament problem. American Review of Mathematics and Statistics, 3(2):7–14, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Dishlieva:2015:OHS

[Dis15] Katya G. Dishlieva. Orbital Hausdorff stability of the solutions of differential equations with variable structure and impulses. American Review of Mathematics and Statistics, 3(1): 70–87, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Fineschi:2016:CPL

[Fin16] Franco Fineschi. A class of piecewise linear maps. American Review of Mathematics and Statistics, 4(2):48–62, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Fode:2014:IGF

[Fod14] Adamou M. Fode. Investigation of Godunov flux against Lax Friedrichs’ flux for the RKDG methods on the scalar nonlinear conservation laws using smoothness indicator. American Review of Mathematics and Statistics, 2(2):43–53, December 2014. CO-DEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Fritz:2021:MPS

[Fri21] Daniel C. Fritz. Measuring the perception strength/beliefs of educators’ of school-based strategies using the Daniel C. Fritz Perception Model to see which is most effective in improving student achievement. American Review of Mathematics and Statistics, 9 (1):41–68, 2021. ISSN 2374-2348 (print), 2374-2356 (electronic).

Gillespie:2015:COC

[GM15] P. Gillespie and S. Molchanov. Covariance operators and the central limit theory for “loop” Markov chains. American Review of Mathematics and Statistics, 3(2):1–6, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Gupta:2022:RSM

[GS22a] Shamita Dutta Gupta and Xiaotie She. On the representative scenario method. American Review of Mathematics and Statistics, 10(1):11–16, June 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Gupta-She:2022:BSM

[GS22b] John Gupta-She. Black–Scholes model and profit and loss attribution of options. American Review of Mathematics and Statistics, 10(2):1–8, December 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Garcia-Santillan:2014:RDM

[GSVMEC14] Arturo García-Santillán, Francisco Venegas-Martínez, and Milka Elena Escalera-Chávez. Restructuring debt model with equivalent equations: Theoretical and practical implications. American Review of Mathematics and Statistics, 2(2):91–106, December 2014. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Gupta:2023:GLR

[Gup23] Shamita Dutta Gupta. Generalized linear regression for ordinal analysis. American Review of Mathematics and Statistics, 11(1): 29–35, June 2023. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Haggag:2018:NFR

[Hag18] Magda M. M. Haggag. A new fuzzy regression model by mixing fuzzy and crisp inputs. American Review of Mathematics and Statistics, 6(2):9–25, December 2018. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Helland:2022:DSM

[Hel22] Inge S. Helland. On the diversity and similarity of mathematical models in science. American Review of Mathematics and Statistics, 10(1):1–10, June 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Hoskuldsson:2023:RRP

[Hös23] Agnar Höskuldsson. Ridge regression and PLS regression. American Review of Mathematics and Statistics, 11(1):1–16, June 2023. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Hou:2018:VSC

[Hou18] Xuezhang Hou. A variable structural control for a flexible plate. American Review of Mathematics and Statistics, 6(2):1–8, December 2018. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Torain:2014:DMM

[II14] David S. Torain II. A dimensionless mathematical model. American Review of Mathematics and Statistics, 2(2):01–26, December 2014. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Izabel:2015:DSR

[Iza15] David Izabel. Do[es] the sequences of random numbers follows a log periodic law? Randomness has memory? American Review of Mathematics and Statistics, 3(1):88–105, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Jamil:2015:CDD

[JA15] Zeana Z. Jamil and Buthainah A. A. Ahmed. Convergence and data dependence result for Picard S-iterative scheme using contractive-like operators. American Review of Mathematics and Statistics, 3(2):83–86, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Jimenez:2019:RFT

[JBA19] Sergio Sedano Jiménez, Javier Ferrero Berlanga, and Gabriel Dorantes Argandar. Rule-following in traffic: a case study for Cuernavaca, Morelos (Mexico). American Review of Mathematics and Statistics, 7(2):10–14, December 2019. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Khodadady:2013:TTM

[KG13] Ebrahim Khodadady and Elmira Ghergloo. S-tests and Ctests: Measures of content-based achievement at grade four of high schools. American Review of Mathematics and Statistics, 1(1):01–16, December 2013. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic). URL http://armsnet.info//

journals/arms/Vol_1_No_1_December_2013/1.pdf.

Khan:2015:NRI

[Kha15] Zareen A. Khan. Nonlinear retarded integral inequalities of Gronwall type. American Review of Mathematics and Statistics, 3(2): 15–23, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Kilburn:2019:AEF

[Kil19] Korey Kilburn. Applying Euler’s formula to integrate e^{(aθ)} sin(bθ) and e^{(aθ)} cos(bθ). American Review of Mathematics and Statistics, 7(2):1–2, December 2019. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Kirmaci:2014:NII

[Kir14] Ugur S. Kirmaci. On new integral inequalities with applications. American Review of Mathematics and Statistics, 2(2):107–124, December 2014. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Kouakou:2019:QTS

[KN19] Rémi K. Kouakou and Firmin K. N’gohissé. Quenching time for some semilinear equations with a potential. American Review of Mathematics and Statistics, 7(2):3–9, December 2019. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Kumar:2016:MDC

[Kum16] Devendra Kumar. Multi dimensional continuous shearlet transform and multi scale wavelet shrinkage with integro differential equations. American Review of Mathematics and Statistics, 4(2): 39–47, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

LaBute:2016:ATV

[La 16] Gerry La Bute. An alternate test for variances. American Review of Mathematics and Statistics, 4(1):11–16, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Li:2018:CST

[LE18] Hong Li and Wayne M. Eby. A comparative study of TELBS robust linear regression. American Review of Mathematics and Statistics, 6(1):1–16, June 2018. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Li:2016:FSF

[Li16] Xianping Li. Fourier series for functions defined on arbitrary limited intervals with polynomial expansion. American Review of Mathematics and Statistics, 4(2):10–17, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Linhares:2017:DBE

[Lin17] Raquel Romes Linhares. Discriminating between explosions and earthquakes by the SDFA method. American Review of Mathematics and Statistics, 5(2):1–5, December 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Linhares:2018:ADC

[Lin18] Raquel Romes Linhares. Arrhythmia detection coefficient based on wavelet shrinkage. American Review of Mathematics and Statistics, 6(2):26–29, December 2018. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Marshall:2015:HTS

[MFH15] David D. Marshall, Brandi N. Falley, and Mark S. Hamner. Hypothesis testing and statistical confidence: an overdue observation on the efficacy of a hypothesis test. American Review of Mathematics and Statistics, 3(1):22–26, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Min:2016:COP

[MHK^{+}16] Kyongyob Min, Keita Hosoi, Yoshinori Kinoshita, Satoshi Hara, Tetsuo Takata, and Ayako Hara. Chronic obstructive pulmonary disease is a disease of multiple heterogeneous conditions changing over time: a Bayesian probability model-based meta-analysis of the ECLIPSE data. American Review of Mathematics and Statistics, 4(2):74–81, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Marshall:2017:CSD

[ML17] David Marshall and Keston Lindsay. Comments and suggestions on defining and using P -values and null hypothesis tests. American Review of Mathematics and Statistics, 5(1):1–6, June 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Motya:2022:PHA

[MMM22] Noureddine Motya, Hakima Mouanis, and Abdelhadi Moutassim. On pre-Hilbert algebras containing a nonzero central idempotent f such that ||fa|| = ||a|| and ||a^{2}|| ≤ ||a||^{2}. American Review of Mathematics and Statistics, 10(1):5–11, June 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Maya:2016:DSD

[MN16] P. Maya and T. Nicholas. Duplication and switching of divisor cordial graphs. American Review of Mathematics and Statistics, 4(2):18–29, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Naidenov:2022:SNR

[Nai22] Nikola Naidenov. Some new representations of the Binet’s function involving Euler sums. American Review of Mathematics and Statistics, 10(1):34–49, June 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Nass:2016:LSI

[NF16] Aminu M. Nass and E. Fredericks. Lie symmetry of Itô stochastic differential equation driven by Poisson process. American Review of Mathematics and Statistics, 4(1):17–30, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Nmah:2022:PTR

[Nma22] Benedict Nmah. A polynomial-time redundancy allocation model. American Review of Mathematics and Statistics, 10(2): 9–25, December 2022. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Nnanwa:2016:POF

[NUE16] C. P. Nnanwa, T. C. Urama, and P. O. Ezepue. Portfolio optimization of financial services stocks in the Nigerian Stock Exchange. American Review of Mathematics and Statistics, 4(2): 1–9, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Otunuga:2014:PSB

[OKL14] Olusegun M. Otunuga, Basant Karna, and Bonita Lawrence. Positive solutions of boundary value dynamic equations. American Review of Mathematics and Statistics, 2(2):55–89, December 2014. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Olapade:2015:DFR

[Ola15] A. K. Olapade. On the distribution functions of the range and the quasi-range of the negatively skewed extended generalized logistic distribution. American Review of Mathematics and Statistics, 3(2):31–38, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

r:2015:SNL

[r15] Haşim Çayır. Some notes on lifts of almost paracontact structures. American Review of Mathematics and Statistics, 3(1):52– 60, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Redmond:2014:MMM

[RCK14] S. P. Redmond, Patricia S. Costello, and Lisa W. Kay. The modern mathematics Master’s degree: a survey of U.S. programs. American Review of Mathematics and Statistics, 2(2):27–42, December 2014. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Redmond:2016:MMM

[RCK16] S. P. Redmond, Patricia S. Costello, and Lisa W. Kay. The modern mathematics Master’s degree II: a survey of U.S. programs. American Review of Mathematics and Statistics, 4(1):1–10, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Rehman:2016:BAI

[Reh16] Najma Abdul Rehman. Biharmonic anti-invariant sub manifolds in kenmotsu space forms. American Review of Mathematics and Statistics, 4(1):73–79, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Rizziolli:2016:DMG

[Riz16] E. C. Rizziolli. On the discriminates of map germs from C^{n+1} to C^{n}, n = 3: 4. American Review of Mathematics and Statistics, 4(1):50–60, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Ruktametakul:2023:ATS

[RRY23] Sarisa Ruktametakul, Sittipong Ruktametakul, and Pornpis Yimprayoon. Application of time series to daily highest and lowest temperatures of Bangkok in Thailand. American Review of Mathematics and Statistics, 11(1):17–21, June 2023. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Sajid:2016:SVO

[Saj16] Dr. Mohammad Sajid. Singular values of one parameter family of function e^{z} + 1/2z. American Review of Mathematics and Statistics, 4(1):45–49, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Sakhabakhsh:2013:EEE

[Sak13] Leila Sakhabakhsh. An evaluation of errors in energy forecasts by the SARFIMA model. American Review of Mathematics and Statistics, 1(1):17–28, December 2013. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic). URL http://armsnet.

info//journals/arms/Vol_1_No_1_December_2013/2.pdf.

Somali:2016:IDC

[SB16] Sennur Somali and Ayse Beler. Iterated defect correction with Bsplines for a class of strongly nonlinear two-point boundary value problems. American Review of Mathematics and Statistics, 4(1): 31–44, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Subedi:2019:ICI

[SI19] Rishi Raj Subedi and James Issos. An improved confidence interval. American Review of Mathematics and Statistics, 7(1):8–11, June 2019. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Sadhasivam:2017:OSF

[SK17] V. Sadhasivam and J. Kavitha. On the oscillation of solutions of fractional vector partial differential equations with deviating arguments. American Review of Mathematics and Statistics, 5 (1):45–57, June 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Song:2015:SAT

[SSJ15] Chwan-Chin Song, Yan-Teng Shih, and Jann-Huei Jinn. A statistical analysis of testlets — a parametric approach. American Review of Mathematics and Statistics, 3(1):34–51, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Tashpulatov:2021:SES

[Tas21] S. M. Tashpulatov. Structure of essential spectra and discrete spectrum of the energy operator of five electron systems in the Hubbard model. Sextet and quartet states. American Review of Mathematics and Statistics, 9(1):12–48, June 2021. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Tuncer:2015:CAS

[TP15] Gökçe Tuncer and Vilda Purutçuoğlu. Comparative assessment of simulation tools for biochemical networks. American Review of Mathematics and Statistics, 3(2):69–82, December 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Tsuchida:2018:LMD

[Tsu18] Shu Tsuchida. Logarithmic mean and the difference calculus. American Review of Mathematics and Statistics, 6(1):17–36, June 2018. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Terzieva:2015:MBP

[TT15] A. Terzieva and G. Terziev. Multitype branching process as model for phytoplankton population and chlorophyll — a contained therein. American Review of Mathematics and Statistics, 3(1):106–111, June 2015. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Udrite:2016:ORA

[UAR16] Constantin Udrişte, Saad Abbas Abed, and Ali Sapeeh Rasheed. Optimal reliability allocation. American Review of Mathematics and Statistics, 4(2):82–91, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Wan:2017:NCS

[WC17] Yong Wan and Weizhi Chen. Normal CR-submanifolds of a quasi Kaehlerian manifold. American Review of Mathematics and Statistics, 5(1):58–67, June 2017. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Wang:2016:LPG

[WY16] Yu Wang and Yong Yu. On lattice properties of GV-semi groups. American Review of Mathematics and Statistics, 4(1): 80–88, June 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).

Wang:2016:APL

[xWX16] Ji xia Wang and Yuewen Xiao. Asymptotic properties of local composite quantile regression estimation for time-varying diffusion model. American Review of Mathematics and Statistics, 4(2):30–38, December 2016. CODEN ???? ISSN 2374-2348 (print), 2374-2356 (electronic).